Twice the larger number is three times the smaller, and their sum is 20. Find the numbers.
Answer by ikdeep(226)

let two numbers be x and y
as per 1st condition;
2x = 3y
or x = 3y/2......(1)
as per 2nd condition
x+y=20......(2)
put the value of 'x' from (1) into (2) and we get
(3y/2) + y = 20
solve the equation and we get,
y = 8
put the value of 'y' into 2 and we get
x=12
hope this will help you